# Probabilistic metaplasticity for continual learning with memristors

### v2

@@ Results
Figure 4a shows the energy consumption per sample during the parameter update phase, and Figure 4b shows the fraction of energy required for computations and memory accesses for the split-MNIST task. We observe that probabilistic metaplasticity leads to a significant reduction in energy. The reason behind this is mainly twofold. First, metaplasticity with gradient accumulation follows the eRBP criteria to assess a weight’s eligibility to update, while probabilistic metaplasticity adds an additional condition of error threshold to update weights. This leads to much fewer weights eligible for update in the latter (up to three orders of magnitude reduction on average), which consequently leads to less frequent execution of the operations needed for weight update. Second, metaplasticity with gradient accumulation needs to compute and store the gradients for the weights eligible for update. This involves three memory accesses for reading the metaplastic coefficient, the current weight gradient, and storing the updated weight gradient to high-precision memory. On the other hand, probabilistic metaplasticity only needs to read the metaplastic coefficient for eligible weights. This leads to 3.3$\sim$ 3.3$\times$ 3.3 × more energy consumption due to memory access for the gradient accumulation approach compared to the probabilistic approach each time the weight update criteria are satisfied. Gradient computation also requires multiplication of the metaplasticity function with the gradient, which consumes more energy compared to the computations required in the probabilistic approach. In summary, gradient accumulation leads to a higher fraction of weights eligible for update, and each eligible weight requires a higher number of memory accesses and computations for gradient computation, leading to higher energy consumption. When the metaplasticity function is complex, such as the exponential function shown in Figure 1c, the gradient accumulation approach requires up to two orders of magnitude higher energy consumption per sample compared to the probabilistic approach. With a bilinear approximation of the exponential function, the energy for the gradient accumulation approach reduces. However, we still observe 60$\sim$ 60$\times$ 60 × reduction in energy consumption per sample with probabilistic metaplasticity.
Figure 5: Effect of individual and shared metaplastic coefficients on energy and stability-plasticity a. Energy consumption during the parameter update phase of the split-MNIST and split-Fashion MNIST tasks for probabilistic metaplasticity with individual and shared metaplastic coefficients. b. The breakdown of energy for computations, SRAM read/write, and memristor read/write for the split-MNIST task. c. Distribution of the product of metaplastic coefficient $m$ and weight magnitude $|w|$ of the hidden layer of the network trained on the split-MNIST task. We see that the shared $m$ leads to lower fraction of weights with low $|mw|$ value. d. The individual task accuracies in the split-MNIST benchmark after sequential training.
@@
Energy analysis shows that SRAM read/write dominates energy consumption in both continual learning approaches. In the probabilistic metaplasticity model, most of the SRAM read/write operations account for updating the metaplastic coefficients assigned to the weights. Here, we optimize the memory overhead by sharing metaplastic coefficients among the weights connected to the same post-synaptic neuron. The metaplastic coefficient associated with the weights connected to the $j^{\text{th}}$ post-synaptic neuron evolves as -
@@
where $X^{\text{tr}}_{j}$ is the activity trace of the $j^{\text{th}}$ post-synaptic neuron.
@@
We evaluate the probabilistic metaplasticity model with shared metaplastic coefficients on a spiking network with 200 hidden neurons and multi-memristor weights ($n_{\text{mem}}$ = 7). Table 3 shows the evaluation of the benchmark tasks. We see that probabilistic metaplasticity with shared metaplastic coefficients performs better compared to state-of-the-art models with weight consolidation based on neural activity [37, 38]. Furthermore, assigning metaplastic coefficients to neurons instead of weights leads to a substantial decrease in memory overhead and energy consumption during training. Energy analysis during the parameter update phase considering a bilinear metaplasticity function shows that shared parameters lead to up to 35$\sim$ 35$\times$ 35 × reduction in energy consumption per sample (Figure 5a). Figure 5b shows the breakdown of energy due to computations, SRAM and memristor read/writes for the split-MNIST task. Shared metaplastic coefficients reduce the number of memory accesses by up to two orders of magnitude, and due to the smaller size of the memory block, each access requires up to 23$\sim$ 23$\times$ 23 × lower energy. This leads to up to three orders of magnitude reduction in energy required for memory accesses. Sharing metaplastic coefficients also improves the scalability of the probabilistic metaplasticity model, as it reduces the space complexity of the algorithm from $O(n^{2})$ to $O(n)$.
@@
The reduction in energy consumption, however, is accompanied by a minor degradation in performance compared to the probabilistic metaplasticity model with individual metaplastic coefficients. This can be attributed to the limited flexibility of the network to consolidate weights, which leads to a poorer balance of the stability-plasticity dilemma. In the split-MNIST task, shared metaplastic parameters lead to reduced plasticity after learning the five sequential tasks. This can be observed from the update probability of the weights, which is inversely proportional to the product of the metaplastic coefficient and the weight magnitude ($|mw|$). The distribution of $|mw|$ in hidden weights (Figure 5c) shows that individual metaplastic coefficients lead to a higher fraction of weights with $|mw|$ close to zero. Individual task accuracies after sequential training also show that shared parameters result in poorer accuracies in later tasks. Despite the reduction in plasticity, the substantial gain in energy consumption incentivizes shared metaplastic coefficients in extreme-edge applications. Introducing mechanisms to reduce the magnitude of metaplastic coefficients based on neural activity can potentially improve the balance of stability and plasticity [39].
@@ Discussion
Emerging nonvolatile memristor devices offer immense opportunities for energy-efficient continual learning on the edge. However, their limited precision and high variability may constrain their suitability for continual learning solutions. Our research takes inspiration from the computational principles of the brain to devise a continual learning mechanism suited to non-ideal low-precision memory devices. Computations in the brain are characterized by stochastic and low-precision synapses [40, 41]. Research in computational memory models demonstrates that stochastic binary synapses with low-precision metaplastic states can help memory retention [26]. Our work draws a parallel between the low-precision hidden metaplastic states of a cascade metaplastic synapse and the conductance levels of a memristor device with two key distinctions. First, instead of binarized weights, we considered multi-level weights represented by memristors. Second, in addition to the state of the synapse, our approach considers the activity of the adjacent neurons to determine the plasticity of a weight. Previous research showed that such activity-dependent metaplasticity leads to a better balance of the stability-plasticity dilemma [11].
@@
In this work, we demonstrated the efficacy of probabilistic metaplasticity in a spiking network trained with error threshold-based eRBP, where weights were realized with a Hafnium oxide-based memristor. However, the proposed mechanism is not limited to the setting above. Probabilistic metaplasticity naturally extends to other surrogate gradient-based training algorithms for spiking networks where the local error at the neurons can be estimated [42]. It can also be explored in conjunction with gradient accumulation-based training for low-precision weights. In this setting, probabilistic metaplasticity will eliminate the need to multiply metaplastic factors with gradients; however, the high frequency of memory accesses may lead to a lower gain in energy consumption. The proposed mechanism can also be explored with other emerging nonvolatile memory devices. Recent research has explored probabilistic training algorithms for emerging low-precision memory devices in deep neural networks considering learning settings with independent and identically distributed data [43, 44]. Probabilistic metaplasticity can be explored in these networks to alleviate catastrophic forgetting during sequential learning with non-stationary data distribution. It is also promising to enable continual learning in neuromorphic accelerators with low-precision weights [45].
@@
We explored the effect of different weight resolutions by employing multi-memristor weights alongside single-memristor weights. Although we observed improved performance with higher resolution weights, it increases the area footprint of memristor crossbars. Improving the resolution of a single memristor device can lead to better performance with lower area overhead. Recent research has shown memristor devices with up to $\sim$ 6-bit precision [21]. In addition to the resolution of conductance modulation, the nonlinearity and variability in memristor devices can affect the performance of error threshold-based training. Due to these non-ideal characteristics, weight updates vary in response to the same local error depending on the current conductance level of the memristor. To minimize this effect, we restricted weight mapping to only ten low-resistance states of the 1T1R device. This choice avoids the higher variability observed in the high-resistance states along with the substantial difference in conductance between the highest low-resistance and high-resistance levels during weight updates[20]. Extreme non-uniformity in the distribution of conductance levels can potentially degrade the performance of error threshold-based training, which can be explored in future studies.
@@ Methods
We train the multilayer spiking neural network with event-driven random backpropagation (eRBP), which uses surrogate gradients [29]. The input stimuli to the network are encoded as Poisson spike trains, where the spike rate is proportional to the strength of the input. The spike train propagates through the weights and accumulates as the input current to the Leaky Integrate and Fire neurons. For the $j^{\text{th}}$ neuron, the accumulated current can be expressed as -
@@
Here, $w_{ij}$ is the weight value, and $m_{ij}$ is a metaplastic coefficient assigned to a weight that captures the activity of its adjacent neurons. It is updated by computing the trace of the neuron activity as -
@@
All datasets used in this work (MNIST [34] and Fashion-MNIST [35]) are publicly available.
@@ Acknowledgements
The authors declare the following competing interests: The authors declare no competing interests.
@@
  • Figure 1: Probabilistic metaplasticity with error threshold-based training. a. Spiking network trained on error threshold where the network weights are realized with 1T1R memristor crossbar array. The dendritic compartments in the hidden and output layer neurons integrate the error. When the error reaches the threshold $U_{\text{th}}$, the memristor weights are updated to the next higher (negative error) or lower (positive error) conductance level with probability $p_{\text{update}}$. b. Mean and standard deviation of resistance levels versus programming compliance current for the 1T1R memristor device (shown in inset) adopted in this work [20]. c. Update probability of weights for different values of the metaplastic coefficient $m$ and the weights $w$. $m$ is directly proportional to the activity level of adjacent neurons. We see that weights with highly active adjacent neurons (high $m$) and weight magnitude (high $|w|$) lead to low update probability.
@@
  • Figure 2: Probabilistic metaplasticity with error threshold-based training for memristor weights. Evolution of task accuracies as the network is trained sequentially on the split-MNIST tasks. The x-axis shows the latest task the network has learned. We evaluate the performance of task $n$ only after the network encounters it and denote the accuracies as 0 before that. a. With no probabilistic metaplasticity, the network learns the current task well, but forgets the initial tasks after sequentially learning multiple tasks. b. Probabilistic metaplasticity with low activity threshold leads to high rigidity in the network, so it remembers previous tasks but cannot learn the last task. c. High activity threshold can lead to loss in previous task accuracy while the network remains plastic to learn the new tasks. d. Optimized activity threshold balances plasticity and rigidity such that the network maintains high initial task accuracies while maintaining the ability to learn new tasks.
@@
  • Figure 3: Mixed-signal architecture and computational flow for the continual learning mechanisms. We consider a mixed-signal architecture where memristor crossbars carry out the forward pass of the spiking model and the error computation and weight update decisions are carried out by digital circuit modules. The figure shows the computational flow for both the proposed mechanism with probabilistic metaplasticity and the gradient accumulation approach with activity-dependent metaplasticity.
  • Figure 4: Energy consumption during parameter update phase. a. Energy required during the parameter update phase per sample with probabilistic metaplasticity and activity-dependent metaplasticity with gradient accumulation considering exponential (exp) and bilinear (lin) metaplasticity functions. b. Breakdown of energy consumption during parameter update in the split-MNIST task. In the diagram, "Computation" denotes all the operations that do not involve SRAM or memristor read/write.
@@
  • Figure 5: Effect of individual and shared metaplastic coefficients on energy and stability-plasticity a. Energy consumption during the parameter update phase of the split-MNIST and split-Fashion MNIST tasks for probabilistic metaplasticity with individual and shared metaplastic coefficients. b. The breakdown of energy for computations, SRAM read/write, and memristor read/write for the split-MNIST task. c. Distribution of the product of metaplastic coefficient $m$ and weight magnitude $|w|$ of the hidden layer of the network trained on the split-MNIST task. We see that the shared $m$ leads to lower fraction of weights with low $|mw|$ value. d. The individual task accuracies in the split-MNIST benchmark after sequential training.
@@
Model Split-MNIST Split-Fashion MNIST Memory Overhead
Baseline 60.69 % $\pm$ 0.6 75.52% $\pm$ 1.31 0 kB
LwF [46] 71.50% $\pm$ 1.63 71.02% $\pm$ 0.46 788 kB
MAS [47] 66.42% $\pm$ 2.47 68.57% $\pm$ 6.85 1.5 MB
BGD [48] 80.44% $\pm$ 0.45 89.73% $\pm$ 0.88 1.5 MB
SS [49] 82.90% $\pm$ 0.01 91.98% $\pm$ 0.12 788 kB
TACOS [11] 82.56% $\pm$ 1.12 93.22% $\pm$ 0.22 1.182 MB
Probabilistic Metaplasticity
$n_{\text{mem}}=1$ 81.14% $\pm$ 1.54 92.56% $\pm$ 1.49 314.4 kB
$n_{\text{mem}}=2$ 82.67% $\pm$ 1.21 92.42% $\pm$ 0.63 314.4 kB
$n_{\text{mem}}=7$ 83.69% $\pm$ 0.78 93.23% $\pm$ 0.14 314.4 kB
Gradient Accumulation w/
metaplasticity ($n_{\text{mem}}=7$) 83.18% $\pm$ 0.42 92.33% $\pm$ 0.27 943.2 kB
@@
Model Split-MNIST Split-FMNIST Memory Overhead
Probabilistic Metaplasticity 83.70 $\pm$ 0.78 93.23 $\pm$ 0.14 314.4 kB
(Individual $m$)
Probabilistic Metaplasticity 82.07 $\pm$ 0.84 92.72 $\pm$ 0.30 $\sim$ 0.5 kB
(Shared $m$)
NAI [37] 68.35 $\pm$ 1.34 68.82 $\pm$ 1.15 3.2 kB
NEO [38] 78.14 $\pm$ 2.23 86.82 $\pm$ 0.60 $\sim$ 1 MB

### v5

@@ Results
Figure 5a shows the energy consumption per sample during the parameter update phase and its breakdown. We highlight the energy consumption due to memory access with memristor and SRAM read/write and subsume the energy required for the rest of the operations (post-synaptic current comparison, random number generation, error computation, etc.) in the "Other operations" category. We observe that probabilistic metaplasticity leads to a significant reduction in energy. The reason behind this is mainly twofold. First, the operations required for updating weights with probabilistic metaplasticity require lower energy compared to the gradient accumulation-based approach. Each time a weight is eligible for update, activity-dependent metaplasticity with gradient accumulation requires computing the gradient and the metaplasticity function. The latter requires reading the metaplasticity coefficient from SRAM and reading memristor weights. The computed gradient is multiplied by the output of the metaplasticity function and accumulated, which requires an SRAM read and write. On the other hand, probabilistic metaplasticity avoids memory access due to gradient accumulation. It also updates weights based on the comparison of a random number with the update probability (computed with the metaplasticity function), which requires much less energy compared to multiplication. Second, in the gradient accumulation-based approach, the weights are updated based on the criteria set in eRBP, while in probabilistic metaplasticity, the eligibility for weight update includes an additional condition of the error threshold. This leads to much fewer weights eligible for update in the latter (up to three orders of magnitude reduction on average), which consequently leads to less frequent execution of the operations needed for weight update. In summary, gradient accumulation leads to a higher fraction of weights eligible for update, and each eligible weight requires a higher number of memory accesses and computations for gradient computation, leading to higher energy consumption. We observe up to 60$\sim$ 60$\times$ 60 × reduction in total energy during parameter updates with probabilistic metaplasticity compared to activity-dependent metaplasticity with gradient accumulation.
@@
Energy analysis shows that SRAM read/write dominates energy consumption in both continual learning approaches. In the probabilistic metaplasticity model, the SRAM read/write operations account for reading and updating the metaplasticity coefficients assigned to the weights. Since a metaplasticity coefficient is assigned to each weight, the area overhead due to this memory can be substantial. Here, we optimize this memory overhead by sharing metaplasticity coefficients among weights. We explore sharing the metaplasticity coefficients based on neural connectivity where weights connected to the same post-synaptic neuron share a metaplasticity coefficient (neuron-shared $m$). The metaplasticity coefficient associated with the weights connected to the $j^{\text{th}}$ post-synaptic neuron evolves as -
@@
Figure 5: Energy consumption and stability-plasticity tradeoff. a. Energy consumption per sample and its breakdown during the parameter update phase for activity-dependent metaplasticity with gradient accumulation and probabilistic metaplasticity with individual and shared metaplasicity coefficients. b. Distribution of the product of metaplasticity coefficient $m$ and weight magnitude $|w|$ in the hidden layer with probabilistic metaplasticity. We see that as more weights share $m$, lower fraction of weights assume low $|mw|$ value. This indicates reduced plasticity in the network. c. Mean accuracy across tasks vs. the energy consumption per sample for the split-MNIST task. We observe a tradeoff between continual learning performance and energy consumption as the metaplasticity coefficients are shared across weights.
where $X^{\text{tr}}_{j}$ is the activity trace of the $j^{\text{th}}$ post-synaptic neuron. This configuration reduces the memory overhead substantially since only one coefficient is required per post-synaptic neuron in a layer. To further optimize the memory overhead, we also consider assigning all weights in the same layer the same metaplasticity coefficient (layer-shared $m$). In this configuration, the metaplasticity coefficient is updated based on the average activity trace of all post-synaptic neurons in the layer -
$m(t+1)=\begin{cases}m(t)+{\Delta}m,&\textrm{\leavevmode\nobreak\ if}\>\frac{(\sum_{j=1}^{n}X^{\text{tr}}_{j})}{n}\geq m^{\text{layer}}_{\text{th}}\\ m(t),&\textrm{\leavevmode\nobreak\ otherwise}\end{cases},$ (5)
where $n$ is the number of post-synaptic neurons in a layer.
@@
We evaluate the probabilistic metaplasticity model with shared metaplasticity coefficients on a spiking network with 200 hidden neurons and multi-memristor weights ($n_{\text{mem}}$ = 7). Table 4 lists the evaluation results on the benchmark tasks. We see that probabilistic metaplasticity with neuron-shared and layer-shared $m$ performs on par or better compared to state-of-the-art models with weight consolidation based on neural activity [43, 42]. Furthermore, sharing metaplasticity coefficients leads to a substantial decrease in memory overhead and energy consumption during training. We model the metaplasticity coefficients neuron-shared and layer-shared $m$ with on-chip SRAM and registers, respectively, during energy analysis. Figure 5a shows, in both neuron-shared and layer-shared configurations, energy consumption is no longer dominated by SRAM read/write, rather most of the energy is spent in operations such as assessing the weight update criteria and memristor read/write. Hence, although layer-shared $m$ drastically reduces memory overhead and leads to three orders of magnitude reduction in energy dissipation to read/write the metaplasticity coefficients compared to neuron-shared $m$, in both cases we observe 18$\sim$ 18$\times$ 18 × reduction in total energy consumption during the parameter update phase.
@@
The reduction in energy consumption, however, is accompanied by a degradation in performance compared to the probabilistic metaplasticity model with individual metaplasticity coefficients. This can be attributed to the limited flexibility of the network to consolidate weights, which leads to a poorer balance of the stability-plasticity dilemma. In the split-MNIST task, sharing metaplasticity coefficients leads to reduced plasticity after sequential training. This can be observed from the update probability of the weights, which is inversely related to the product of the metaplasticity coefficient and the weight magnitude ($|mw|$). The distribution of $|mw|$ in hidden weights (Figure 5b) shows that shared metaplasticity coefficients lead to a lower fraction of weights with $|mw|$ close to zero, indicating reduced plasticity. This reduction is more pronounced in layer-shared $m$. Individual task accuracies after sequential training also show that shared coefficients result in poorer accuracies in later tasks (Supplementary Table S5). This demonstrates a trade-off in the memory overhead and the network’s ability to balance stability and plasticity.
Despite the reduction in plasticity, the substantial gain in energy consumption incentivizes shared metaplasticity coefficients in extreme-edge applications. The trade-off between memory overhead and the stability-plasticity balance can be further optimized by sharing the metaplasticity coefficients among blocks of weights within a layer. We investigated a modular sharing of metaplasticity coefficients (module-shared $m$), where a block of weights, connecting a post-synaptic neuron to a set of pre-synaptic neurons, shares a metaplasticity coefficient. The pre-synaptic neurons can be selected based on different criteria such as the sign or magnitude of incoming weights. Here, we group pre-synaptic neurons based on proximity where adjacent pre-synaptic neurons are grouped together. This configuration is well-suited for hardware implementation, as the activity of adjacent neurons can be computed locally, minimizing communication overhead. The coefficient $m_{kj}$, assigned to module $k$, consisting of post-synaptic neuron j and $n_{\text{block}}$ number of adjacent pre-synaptic neurons, are updated based on the post-synaptic activity trace and the average of the traces of the pre-synaptic neurons in the module as follows -
$m_{kj}(t+1)=\begin{cases}m_{kj}(t)+{\Delta}m,&\textrm{\leavevmode\nobreak\ if}\>\frac{(\sum_{i=1}^{n_{\text{block}}}X^{\text{tr}}_{i})}{n_{\text{block}}}\geq m^{\text{pre}}_{\text{th}}\&X^{\text{tr}}_{j}\geq m^{\text{post}}_{\text{th}}\\ m_{kj}(t)&\textrm{\leavevmode\nobreak\ otherwise}\end{cases}.$ (6)
We evaluate the performance of probabilistic metaplasticity with module-shared $m$ in the split-MNIST task considering sharing $m$ among 8 and 4 adjacent weights in the hidden and the output layer, respectively. Figure 5c shows the mean accuracy across tasks for different configurations of shared $m$ vs. the energy consumption per sample during the parameter update phase. We see that while the performance of module-shared $m$ is similar to that of neuron-shared $m$, the former shows greater plasticity compared to neuron-shared $m$ (Figure 5b). It also reduces memory overhead by $\sim 87.5\%$ and reduces energy consumption by 4.7$\sim$ 4.7$\times$ 4.7 × during the parameter update phase compared to individual $m$. This configuration can be a promising approach to optimize the trade-off between balance of stability and plasticity and energy efficiency. Introducing mechanisms to reduce the magnitude of metaplasticity coefficients based on neural activity can also potentially improve the balance of stability and plasticity [44].
@@ Discussion
Emerging nonvolatile memristor devices offer immense opportunities for energy-efficient continual learning on the edge. However, their limited precision and high variability may constrain their suitability for continual learning solutions. Our research takes inspiration from the computational principles of the brain to devise a continual learning mechanism suited to non-ideal low-precision memory devices. Computations in the brain are characterized by stochastic and low-precision synapses [45, 46]. Research in computational memory models demonstrates that stochastic binary synapses with low-precision metaplastic states can help memory retention [26]. Our work draws a parallel between the low-precision hidden metaplastic states of a cascade metaplastic synapse and the conductance levels of a memristor device. Abstracting this concept, we interpret the plasticity of weights as their probability of update rather than the magnitude of the update. This interpretation distinguishes the proposed work from previous continual learning models with metaplasticity. Previous work has also explored probabilistic updates in memristor weights considering learning settings with independent and identically distributed (i.i.d.) data [47, 48]. In a continual learning setting, where the distribution of data changes, such mechanisms are prone to catastrophic forgetting. In contrast, probabilistic metaplasticity effectively alleviates catastrophic forgetting by dynamically adjusting the update probability of weights based on their importance.
@@
In this work, we demonstrated the efficacy of probabilistic metaplasticity in a spiking network trained with error threshold-based eRBP, where weights were realized with a Hafnium oxide-based memristor. However, the proposed mechanism is not limited to the setting above and can be explored in the context of different network architectures and memory devices. The proposed work naturally extends to other surrogate gradient-based training algorithms for spiking networks, where the local error at the neurons can be estimated [49]. Additionally, it can be applied to deep neural networks with low-precision weights updated using stochastic rounding [50]. Incorporating probabilistic metaplasticity with gradient accumulation-based training is also possible. In this setting, probabilistic metaplasticity will eliminate the need to multiply metaplasticity factors with gradients. However, the high frequency of memory access may lead to a lower gain in energy consumption. The proposed mechanism also extends to networks that realize low-precision weights with other memory devices. It shows promise to enable continual learning in neuromorphic accelerators with low-precision weights [51]. Recent research has explored probabilistic training algorithms for emerging low-precision memory devices in deep neural networks considering learning settings with independent and identically distributed data [47, 48]. Probabilistic metaplasticity can be explored in these networks to alleviate catastrophic forgetting during sequential learning with non-stationary data distribution.
@@
We explored the effect of different weight resolutions by employing multi-memristor weights alongside single-memristor weights. Although we observed improved performance with higher resolution weights, it increases the area footprint of memristor crossbars. Improving the resolution of a single memristor device can lead to better performance with lower area overhead. Recent research has shown memristor devices with up to $\sim$ 6-bit precision [21]. In addition to the resolution of conductance modulation, the nonlinearity and variability in memristor devices can affect the performance of error threshold-based training. Due to these non-ideal characteristics, weight updates vary in response to the same local error depending on the current conductance level of the memristor. To minimize this effect, we restricted weight mapping to only ten low-resistance levels of the 1T1R device. This choice avoids the higher variability observed in the high-resistance levels along with the substantial difference in conductance between the highest low-resistance and high-resistance levels during weight updates[20]. However, it should be noted that choosing low-resistance levels increases parasitic voltage drop and sneak currents, which can also introduce nonlinearity and degrade the network performance. Similarly, extreme non-uniformity in the distribution of conductance levels can potentially degrade the performance of error threshold-based training. Future studies incorporating optimized circuits and system demonstrations of the proposed mechanism can further investigate the effect of these non-idealities.
This work limits the exploration of probabilistic metaplasticity in fully-connected spiking networks. Future research can expand the proposed mechanism to larger networks and tackle more complex continual learning tasks with longer sequences. It would be interesting to extend the proposed mechanism to convolutional spiking networks where weights are shared among inputs. We identified a trade-off between memory overhead and the network’s ability to balance stability and plasticity when metaplasticity coefficients are shared among network weights. Further theoretical analysis could offer valuable insights and pave the way for more efficient continual learning solutions. Further investigations into the contrasting learning dynamics when training the network based on error accumulated at neurons instead of gradients can also lead to a deeper understanding of the underlying mechanisms. Our empirical observations on continual learning tasks did not show any significant differences in performance (Supplementary Table S6). However, we note that gradient accumulation may allow for more gradual weight changes over time as the gradients can be accumulated over samples. During error threshold-based training, integrating error across samples may lead to incorrect attribution, especially when weights are updated based on immediate input activity. For settings that require gradual updates, tracing input activity over time or using low-precision counters to track smaller changes may prove beneficial. Error-triggered training is promising for resource-constrained platforms as it reduces memory overhead and merits robust study under various conditions.
@@ Methods
We train the multilayer spiking neural network with event-driven random backpropagation (eRBP) combined with error threshold, which uses surrogate gradients [29]. This section describes the eRBP training rule. The input stimuli to the network are encoded as Poisson spike trains, where the spike rate is proportional to the strength of the input. The spike train propagates through the weights and accumulates as the input current to the Leaky Integrate and Fire neurons. For the $j^{\text{th}}$ neuron, the accumulated current can be expressed as -
@@
Here, $w_{ij}$ is the weight value, and $m_{ij}$ is a metaplasticity coefficient assigned to a weight that captures the activity of its adjacent neurons. It is updated by computing the trace of the neuron activity as -
@@
All datasets used in this work (MNIST [34], Fashion-MNIST [35] and CIFAR-10 [41]) are publicly available.
@@ Acknowledgements
The authors declare no competing interests.
@@
  • Figure 1: Probabilistic metaplasticity with error threshold-based training. a. Spiking network trained on error threshold where the network weights are realized with 1T1R memristor crossbar array. The dendritic compartments in the hidden and output layer neurons integrate the error. When the error reaches the threshold $U_{\text{th}}$, the memristor weights are updated to the next higher (negative error) or lower (positive error) conductance level with probability $p_{\text{update}}$. b. Mean and standard deviation of resistance levels versus programming compliance current for the 1T1R memristor device (shown in inset) adopted in this work [20]. c. Update probability of weights for different values of the metaplasticity coefficient $m$ and the weights $w$. $m$ is positively associated with the activity level of adjacent neurons. We see that weights with highly active adjacent neurons (high $m$) and weight magnitude (high $|w|$) lead to low update probability.
@@
  • Figure 2: Evolution of task accuracies with sequential training on the split-MNIST benchmark. The x-axis shows the latest task the network has learned. We evaluate the performance of task $n$ only after the network encounters it and denote the accuracies as 0 before that. a. With no probabilistic metaplasticity, the network learns the current task well, but forgets the initial tasks after sequentially learning multiple tasks. b. Probabilistic metaplasticity with low activity threshold leads to high rigidity in the network, so it remembers previous tasks but cannot learn the last task. c. High activity threshold can lead to loss in previous task accuracy while the network remains plastic to learn the new tasks. d. Optimized activity threshold balances plasticity and rigidity such that the network maintains high initial task accuracies while maintaining the ability to learn new tasks.
  • Figure 3: Effect of number of training samples on continual learning. The mean accuracy across tasks shows degradation as the average number of samples per task reduces.
@@
  • Figure 4: Mixed-signal architecture and computational flow for the continual learning mechanisms. We consider a mixed-signal architecture where memristor crossbars carry out the forward pass of the spiking model and the error computation, arithmetic and logical operations are carried out in the digital domain. The figure shows the computational flow for continual learning with probabilistic metaplasticity (left) and activity-dependent metaplasticity with gradient accumulation (right).
@@
  • Energy consumption and stability-plasticity tradeoff. a. Energy consumption per sample and its breakdown during the parameter update phase for activity-dependent metaplasticity with gradient accumulation and probabilistic metaplasticity with individual and shared metaplasicity coefficients. b. Distribution of the product of metaplasticity coefficient $m$ and weight magnitude $|w|$ in the hidden layer with probabilistic metaplasticity. We see that as more weights share $m$, lower fraction of weights assume low $|mw|$ value. This indicates reduced plasticity in the network. c. Mean accuracy across tasks vs. the energy consumption per sample for the split-MNIST task. We observe a tradeoff between continual learning performance and energy consumption as the metaplasticity coefficients are shared across weights.
@@
Model Split-MNIST Split-Fashion MNIST Normalized Memory Overhead Baseline 60.69 % $\pm$ 0.6 75.52% $\pm$ 1.31 0 kB LwF [39] 71.50% $\pm$ 1.63 71.02% $\pm$ 0.46 628.8 kB MAS [36] 66.42% $\pm$ 2.47 68.57% $\pm$ 6.85 $\sim$ 1.2 MB BGD [38] 80.44% $\pm$ 0.45 89.73% $\pm$ 0.88 $\sim$ 1.2 MB SS [37] 82.90% $\pm$ 0.01 91.98% $\pm$ 0.12 628.8 kB TACOS [11] 82.56% $\pm$ 1.12 93.22% $\pm$ 0.22 $\sim$ 0.9 MB Probabilistic Metaplasticity $n_{\text{mem}}=1$ 81.14% $\pm$ 1.54 92.56% $\pm$ 1.49 $\sim$ 0.3 MB $n_{\text{mem}}=2$ 82.67% $\pm$ 1.21 92.42% $\pm$ 0.63 $\sim$ 0.3 MB $n_{\text{mem}}=7$ 83.69% $\pm$ 0.78 93.23% $\pm$ 0.14 $\sim$ 0.3 MB Activity-dependent Metaplasticity w/ Gradient Accumulation ($n_{\text{mem}}=7$) 83.18% $\pm$ 0.42 92.33% $\pm$ 0.27 $\sim$ 0.9 MB
@@
Table 3: Evaluation of a spiking neural network with probabilistic metaplasticity and TACOS on the split-CIFAR-10 task in the domain-IL scenario. Probabilistic metaplasticity is evaluated considering multi-memristor weights $n_{mem}$ = 7 and TACOS considers full-precision weights. The table lists the mean and standard deviation of the individual task accuracies and the mean accuracy across tasks over 5 runs after sequential training.
Task Baseline Probabilistic Metaplasticity TACOS[11]
Class 0,1 86.64 $\pm$ 1.86 93.52 $\pm$ 1.36 90.84 $\pm$ 0.36
Class 2,3 59.81 $\pm$ 3.69 73.11 $\pm$ 4.80 71.27 $\pm$ 1.08
Class 4,5 63.93 $\pm$ 3.08 80.87 $\pm$ 4.07 80.01 $\pm$0.93
Class 6,7 82.03 $\pm$ 5.75 84.90 $\pm$ 1.61 88.83 $\pm$ 0.57
Class 8,9 96.96 $\pm$ 0.80 94.71 $\pm$ 0.21 97.64 $\pm$ 0.23
Mean 77.87 $\pm$ 2.77 85.42 $\pm$ 2.11 85.72 $\pm$ 0.55
@@
Model Split-MNIST Split-FMNIST Normalized Memory
Overhead
Probabilistic Metaplasticity
Individual m 83.70 $\pm$ 0.78 93.23 $\pm$ 0.14 $\sim$ 0.3 MB
Neuron-shared m 82.07 $\pm$ 0.84 92.72 $\pm$ 0.30 $\sim$ 0.4 kB
Layer-shared m 79.74 $\pm$ 1.21 92.01 $\pm$ 0.49 4 Bytes
NAI [43] 68.35 $\pm$ 1.34 68.82 $\pm$ 1.15 $\sim$ 0.8 kB
NEO [42] 78.14 $\pm$ 2.23 86.82 $\pm$ 0.60 $\sim$ 0.3 MB